OpenStudy (anonymous):

Dorothy ran 6 times around a circular track that has a diameter of 47 m. Approximately how far did she run? Use pie = 3.14 and round your answer to the nearest meter.

OpenStudy (anonymous):

You want to use the circumference to make your life easier... The circumference of the track= π * diameter or π * d = (3.14) * (47 meters) = 148.52 meters Dorothy ran around the track 6 times so Distance Dorothy ran = 148.52 meters * 6 = 891.12 meters Dorothy ran about 891 meters. PS. Its pi not pie, hehe.

OpenStudy (anonymous):

nice